Voltage across / resistor = 1 R = 0.58 A × 7 = 4.06 V


1
Expert's answer
2021-05-06T16:52:17-0400



Voltage across "7\ \Omega" register:

For the given circuit

The resultant resistance is

"R_{eq}=7+5||10\\R_{eq}=7+\dfrac{10\times5}{10+5}\\R_{eq}=\dfrac{155}{15}=10.33\ \Omega"


Total Current is

"I=\dfrac{V}{R_{eq}}\\I=\dfrac{6}{10.33}=0.58\ A"


So, Voltage across "7\ \Omega" register is

"V_7=IR_7=0.58\times 7=4.06\ V"


Hence, Voltage across "7\ \Omega" register is 4.06 V